# A Survey of Quantized Graph Representation Learning: Connecting Graph Structures with Large Language Models

### v1

@@ 3.4 KG Quantization
Recently, several KG quantization methods have been proposed for effective embedding, responding to the increasing demand for larger KGs. Unlike the techniques employed in the HOG or HEG setting, these methods tend to leverage an unsupervised paradigm for quantization. NodePiece Galkin et al. (2022), EARL Chen et al. (2023), and random entity quantization (RandomEQ for short) Li et al. (2023a) are all in such a framework. They first select a number of entities as anchors and then utilize the structural statistical strategy to match them for each entity. Specifically, NodePiece employs metrics such as Personalized PageRank Page (1999) and node degree to select certain anchor entities and subsequently view top- closest anchors as entity codes. EARL and RandomEQ sample 10% entities as anchors and then assign soft weights to each anchor according to the similarity between connected relation sets. Given the limitations of these methods in thoroughly capturing the structure and semantics of KGs, SSQR presents a self-supervised approach that employs learned discrete codes to reconstruct structures and imply semantic text, offering more holistic modeling with only 16 quantized codes.
@@ 3.5 Application Scenarios
General Graph Tasks.
Based on the learned quantized codes, many general graph tasks can be addressed, for example, node classification, link prediction, graph classification, graph regression, and graph generation. In addition, SNEQ He et al. (2020) can be employed for node recommendation that ranks all nodes based on a specific distance metric and suggests the nearest node. d-SNEQ He et al. (2023) is further utilized for path prediction that predicts the path (e.g., the shortest path) between two nodes.
@@
Molecular Tasks & AI for Science.
Recently, QGR methods have been widely used for molecular tasks and AI for science scenarios, including molecular property prediction (classification and regression), molecule-text prediction, and protein & RNA reconstruction. Specifically, DGAE Boget et al. (2024a) conduct the graph generation. It first iteratively samples discrete codes and then generates the graph structures by the pre-trained decoder, which can be used in the generation for molecular graphs and has the potential for drug design, material design and protein design. LLPS Gaujac et al. (2024) shows the potential for the reconstruction of protein sequences by training a de novo generative model for protein structures using a vanilla decoder-only Transformer model. Bio2Token Liu et al. (2024a) conducts efficient representation of large 3D molecular structures with high fidelity for molecules, proteins, and RNA, holding the potential for the design of biomolecules and biomolecular complexes. UniMoT Zhang et al. (2024) unifies the molecule-text applications, including molecular classification & regression, molecule captioning, molecule-text retrieval, and caption-guided molecule generation. It demonstrates the impressive and comprehensive potentials that arise from the combination of QGR and LLMs.
@@ 4 Code Dependence Learning
where network is with parameter . is the length of the code sequence c. DGAE Boget et al. (2024a) splits the latent representation of each node into parts. So after the quantization, each graph can be represented as code permutation. To learn the dependence of it, the 2D Transformer Vaswani et al. (2017) is introduced to auto-regressively generate the codes, i.e., the joint probability can be .
Moreover, GLAD Boget et al. (2024b) implements diffusion bridges Liu et al. (2023b) for codes’ dependence, where Brownian motion is utilized as a non-conditional diffusion process defined by a stochastic differential equation (SDE). Based on the learned model bridge, discrete practical codes can be acquired by the iterative denoising process from the random Gaussian noise. Followed by the pre-trained decoder, these codes can be used for the graph generation. Although Dr.E does not directly or explicitly learn the dependence of the codes, it introduces intra-layer and inter-layer residuals to gradually generate the codes, which enhances the representation of sequential information as the newly generated code depends on the previously obtained codes. It can be viewed as an implicit auto-regressive manner.
@@ 5 Integration with LLMs
(a) Node classification, taken from Dr.E Liu et al. (2024b). |
(b) KG triple classification, taken from SSQR Lin et al. (2025a). |
@@
Dr.E implements token-level alignment between GNNs and LLMs. Based on the quantized codes through intra-layer and inter-layer residuals, the preservation of multiple perspectives in each convolution step is guaranteed, fostering reliable information transfer from the surrounding nodes to the central point. Moreover, by utilizing the LLMs’ vocabulary as the codebook, Dr.E can represent each node with real words from different graph hops, as shown in Table 2 (a). By fine-tuning LLaMA-2-7B Touvron et al. (2023) with the designed instruction data, the node label can be easily predicted. Different from the fixed vocabulary in Dr.E, UniMoT Zhang et al. (2024) expands the original LLMs’ vocabulary through the incorporation of learned molecular tokens from the codebook, forming a unified molecule-text vocabulary. Based on the specific instructions for molecular property prediction, molecule captioning, molecule-text retrieval, caption-guided molecule generation, LLaMA2 is fine-tuned with LoRA Hu et al. (2022), thereby bestowing upon it an impressive and comprehensive proficiency in molecule-text applications. Similarly, in KG-related scenarios, SSRQ leverages the learned discrete entity codes as additional features to directly input LLMs, helping to make accurate predictions. Both LLaMA-2-7B and LLaMA3.1-8B Dubey et al. (2024) are utilized for the KG link prediction task and triple classification task. During the fine-tuning phase, the introduction of new tokens necessitates the expansion of the LLM’s tokenizer vocabulary, which is in line with the size of the codebook, e.g., 1024 or 2048. All other elements of the network remain unaltered.
6 Discussions and Future Directions
Beyond the above methods, there are also some methods that does not use the quantized code for the final representation, such as VQ-GNN Ding et al. (2021), VQSynery Wu et al. (2024), and MSPmol Lu et al. (2024). For example, instead of employing discrete motif features as the ultimate representation, MSPmol Lu et al. (2024) utilizes the Vector Quantization (VQ) method on motif nodes within specific layers of the molecular graph. There are also benefits for the representation, which can be viewed as a paradigm of contiguous+quantized with the network, rather than the outputs. They differ from the studies in Table 1, so we do not include them in this survey.
Despite some achievements of the QGR methods, there are also some shortcomings from both methodology and application perspectives. First, the quantization process may result in the loss of some details of the original data, which could pose challenges in areas that require accurate modeling. Second, the quantization process can require complex techniques to ensure effective and efficient optimization, while pinepine’s schema and integration with LLMs add additional complexity to the overall framework. Third, compared with the wide application and success of general LLMs, the research on the integration of QGR and LLMs is still lacking. Thus, the direction of future development lies in the following four aspects.
Choice of Codeword.
@@ 6 Discussions and Future Directions
For QGR learning, one can incorporate certain domain knowledge into the construction of the codebook. This knowledge might encompass semantic concepts, structural formations, and textual language. Such an approach can augment the interpretability of the whole framework, thereby facilitating a better human understanding.
Advanced Quantized Methods.
@@
GQR can adopt the latest advanced quantized methods to enhance the efficiency and efficacy of learning graph codes, which may have already demonstrated success within the domain of computer vision. For instance, the rotation trick Fifty et al. (2024) is a proven method to mitigate the issues of codebook collapse and underutilization in VQ. SimVQ Zhu et al. (2024) simply incorporates a linear transformation into the codebook, yielding substantial improvement.
Unified Graph Foundation Model.
@@
While the GQR methods have indeed made some notable strides, the current research primarily follows a similar approach that tackles various tasks using distinct training targets, resulting in multiple distinct models. This could restrict its applicability, particularly in the age of LLMs. Drawing inspiration from the remarkable success of unified LLMs and multimodal LLMs Zhao et al. (2023); Lin et al. (2025b), the unified graph foundation model could be realized to execute various tasks across diverse graphs through QGR, by learning unified codes for different graphs and then tuning with LLM techniques.
Graph RAG with QGR.
@@
Retrieval-Augmented Generation (RAG) has emerged as a significant focal point for improving the capabilities of LLMs within specific fields. Graph RAG Edge et al. (2024) would be beneficial for downstream tasks in the context of LLMs, where the QGR learning would provide an effective manner to retrieve relevant nodes or subgraphs by calculating the similarity of the codes. It can be easily realized by the statistical metrics of the codes or the numerical calculations among embeddings.

### v2

@@ 3.4 KG Quantization
Recently, several KG quantization methods have been proposed for effective embedding, responding to the increasing demand for larger KGs. Unlike the techniques employed in the HOG or HEG setting, these methods tend to leverage an unsupervised paradigm for quantization. NodePiece Galkin et al. (2022), EARL Chen et al. (2023), and random entity quantization (RandomEQ for short) Li et al. (2023a) are all in such a framework. They first select a number of entities as anchors and then utilize the structural statistical strategy to match them for each entity. Specifically, NodePiece employs metrics such as Personalized PageRank Page (1999) and node degree to select certain anchor entities and subsequently view top- closest anchors as entity codes. EARL and RandomEQ sample 10% entities as anchors and then assign soft weights to each anchor according to the similarity between connected relation sets. Given the limitations of these methods in thoroughly capturing the structure and semantics of KGs, SSQR Lin et al. (2025a) presents a self-supervised approach that employs learned discrete codes to reconstruct structures and imply semantic text, offering more holistic modeling with only 16 quantized codes. Furthermore, ReaLM Guo et al. (2025) introduces an RVQ framework that compresses high-dimensional and continuous KG embeddings (pretrained RotatE Sun et al. (2019) embeddings) into compact and discrete code sequences while effectively preserving rich relational and semantic information. More detailed comparison between SSQR and ReaLM are shown in Appendix §D.
@@ 3.5 Application Scenarios
General Graph Tasks. Based on the learned quantized codes, many general graph tasks can be addressed, for example, node classification, link prediction, graph classification, graph regression, and graph generation. In addition, SNEQ He et al. (2020) can be employed for node recommendation that ranks all nodes based on a specific distance metric and suggests the nearest node. d-SNEQ He et al. (2023) is further utilized for path prediction that predicts the path (e.g., the shortest path) between two nodes.
@@
Molecular Tasks & AI for Science. Recently, QGR methods have been widely used for molecular tasks and AI for science scenarios Chen et al. (2024), including molecular property prediction (classification and regression), molecule-text prediction, and protein & RNA reconstruction. Specifically, DGAE Boget et al. (2024) conduct the graph generation. It first iteratively samples discrete codes and then generates the graph structures by the pre-trained decoder, which can be used in the generation for molecular graphs and has the potential for drug design, material design and protein design. LLPS Donà et al. (2025) shows the potential for the reconstruction of protein sequences by training a de novo generative model for protein structures using a vanilla decoder-only Transformer model. Bio2Token Liu et al. (2024) conducts efficient representation of large 3D molecular structures with high fidelity for molecules, proteins, and RNA, holding the potential for the design of biomolecules and biomolecular complexes. UniMoT Zhang et al. (2024) unifies the molecule-text applications, including molecular classification & regression, molecule captioning, molecule-text retrieval, and caption-guided molecule generation.
@@ 4 Code Dependence Learning
where network is with parameter . is the length of the code sequence c. DGAE Boget et al. (2024) splits the latent representation of each node into parts. So after the quantization, each graph can be represented as code permutation. To learn the dependence of it, the 2D Transformer Vaswani et al. (2017) is introduced to auto-regressively generate the codes, i.e., the joint probability can be .
Moreover, GLAD Nguyen et al. (2025) implements diffusion bridges Liu et al. (2023b) for codes’ dependence, where Brownian motion is utilized as a non-conditional diffusion process defined by a stochastic differential equation (SDE). Based on the learned model bridge, discrete practical codes can be acquired by the iterative denoising process from the random Gaussian noise. Followed by the pre-trained decoder, these codes can be used for the graph generation. Although Dr.E does not directly or explicitly learn the dependence of the codes, it introduces intra-layer and inter-layer residuals to gradually generate the codes, which enhances the representation of sequential information as the newly generated code depends on the previously obtained codes. It can be viewed as an implicit auto-regressive manner.
@@ 5 Integration with LLMs
| (a) Node classification, taken from Dr.E Liu et al. (2025). |
| (b) KG triple classification, from SSQR Lin et al. (2025a). |
@@
Dr.E implements token-level alignment between GNNs and LLMs. Based on the quantized codes through intra-layer and inter-layer residuals, the preservation of multiple perspectives in each convolution step is guaranteed, fostering reliable information transfer from the surrounding nodes to the central point. Moreover, by utilizing the LLMs’ vocabulary as the codebook, Dr.E can represent each node with real words from different graph hops, as shown in Table 2 (a). By fine-tuning LLaMA-2-7B Touvron et al. (2023) with the designed instruction data, the node label can be easily predicted. Different from the fixed vocabulary in Dr.E, UniMoT Zhang et al. (2024) expands the original LLMs’ vocabulary through the incorporation of learned molecular tokens from the codebook, forming a unified molecule-text vocabulary. Based on the specific instructions for molecular property prediction, molecule captioning, molecule-text retrieval, caption-guided molecule generation, LLaMA2 is fine-tuned with LoRA Hu et al. (2022), thereby bestowing upon it an impressive and comprehensive proficiency in molecule-text applications. Similarly, in KG-related scenarios, SSRQ leverages the learned discrete entity codes as additional features to directly input LLMs, helping to make accurate predictions. Both LLaMA2-7B and LLaMA3.1-8B Dubey et al. (2024) are utilized for the KG link prediction task and triple classification task. During the fine-tuning phase, the introduction of new tokens necessitates the expansion of the LLM’s tokenizer vocabulary, which is in line with the size of the codebook, e.g., 1024 or 2048. All other elements of the network remain unaltered. Similarly, ReaLM employs LoRA fine-tuning on Llama3.2 to adapt the model for KG tasks.
6 Future Directions
Based on this systematic review, we identify four key directions for future research.
@@ 6 Future Directions
Choice of Codeword. For QGR learning, one can incorporate certain domain knowledge into the construction of the codebook. This knowledge might encompass semantic concepts, structural formations, and textual language. Such an approach can augment the interpretability of the whole framework, facilitating a better human understanding.
@@
Advanced Quantized Methods. QGR can leverage cutting-edge VQ methods to enhance the effectiveness and efficiency of graph codes, drawing on techniques that have already proven successful in the computer vision domain. For instance, the rotation trick Fifty et al. (2024) is a proven method to mitigate the issues of codebook collapse and underutilization in VQ. SimVQ Zhu et al. (2024) simply incorporates a linear transformation into the codebook, yielding substantial improvement.
@@
Unified Graph Foundation Model. While the QGR methods have indeed made some notable strides, the current research primarily follows a schema that tackles various tasks using distinct training targets, resulting in multiple models. This diversity may restrict its applicability, particularly in the age of LLMs. Drawing inspiration from the remarkable success of unified LLMs and multimodal LLMs Zhao et al. (2023); Lin et al. (2025b), the unified graph foundation model could be realized to execute various tasks across diverse graphs through QGR, by learning unified codes for different graphs and then tuning with LLM techniques Ji et al. (2025).
@@
Graph RAG with QGR. Retrieval-Augmented Generation (RAG) has emerged as a significant focal point for improving the capabilities of LLMs within specific fields. Graph RAG Edge et al. (2024); Luo et al. (2025) would be beneficial for downstream tasks in the context of LLMs, where the QGR learning would provide an effective manner to retrieve relevant nodes or subgraphs by calculating the similarity of the codes. It can be easily realized by the statistical metrics of the codes or the numerical calculations among embeddings.
Limitations
Beyond the above methods, there are also some methods that do not use quantized code for the final representation, such as VQ-GNN Ding et al. (2021), VQSynery Wu et al. (2024), and MSPmol Lu et al. (2024). For example, instead of employing discrete motif features as the ultimate representation, MSPmol Lu et al. (2024) utilizes the Vector Quantization (VQ) method on motif nodes within specific layers of the molecular graph. There are also benefits for the representation, which can be viewed as a paradigm of contiguous+quantized with the network, rather than the outputs. They differ from the studies in Table 1, so we do not include them in this survey.
Despite some achievements of the QGR methods, there are still some limitations from both methodology and application perspectives Li et al. (2025). First, the quantization process may result in the loss of some details of the original data, which could pose challenges in areas that require accurate modeling. Second, the quantization process requires complex techniques to ensure effective and efficient optimization, while pinepine’s schema and integration with LLMs further increase the complexity of the overall framework. Third, compared with the wide application and success of general LLMs, efforts to integrate QGR and LLMs still need to be carried out. Leveraging multi-agent frameworks is a compelling direction Zhang et al. (2026a, b), as it could learn semantic representations to substantially refine the planning and collaborative reasoning capabilities of these systems.
Appendix A Merits of Quantized Graph Representation
Embedding Parameter Efficiency. As large-scale graphs become ubiquitous in reality, there is a corresponding substantial increase in the representation parameters, which necessitates abundant memory usage for storage and substantial computational expenses Galkin et al. (2022). Supposing as the number of nodes and as the dimensions, the required embedding parameters would be . In the quantized settings, the total count of parameters would reduce to , where and denote the length of the codeword set and the number of codes assigned to each node. For example, if there are nodes with 1024 feature dimensions, 1024 codewords, and 8 assigned codes, it would be a 113-fold reduction in required representation parameters.
Explainability and Interpretability. The discrete nature of QGR brings the explainability and interpretability for embeddings and reasoning, as shown in Figure1 (b). Each code would be assigned to the practical item by specific designs for direct interpretability, e.g., Dr. E Liu et al. (2025) introduces a language vocabulary as the code set and each node in graphs can be represented as the permutation of explainable language tokens.
Robustness and Generalization. The discrete tokens of the QGR are more robust and generalizable, which usually utilizes the self-supervised strategies to learn both local-level and high-level graph structures (e.g., edge reconstruction and DGI Velickovic et al. (2019)) as well as semantics (e.g., feature reconstruction). Moreover, the acquired tokens can readily adapt to downstream applications directly or indirectly Yang et al. (2024).
Seamless Integration with NLP Models. The swift advancement in natural language processing (NLP) techniques, particularly LLMs Xu et al. (2024), has sparked an increased interest in employing NLP models to resolve graph tasks. However, the inherent representation gap between the typical graph structure and natural language poses a significant challenge to their seamless and effective integration. Graph quantization, by learning discrete codes that bear resemblance to natural language forms, can facilitate this integration directly and seamlessly Lin et al. (2025a); Liu et al. (2025).
Appendix B Algorithm for RVQ
To present the process more intuitively and clearly, we summarize the entire RVQ procedure in the following algorithm.
Appendix C General Framework of QGR Studies
The general framework of the QGR studies is illustrated in Figure 2, which mainly comprises an encoder , decoder , and quantization process . The encoder is to model the original graph structure into latent space, where MLPs He et al. (2020) and GNNs Xia et al. (2023) are usually utilized. Training objectives of different levels can be utilized. By combining a predictor, multiple applications (including node level, edge level and graph level) can be realized.
Appendix D Comparison between SSQR and ReaLM
From a methodological perspective, SSQR is optimized through structure reconstruction and semantic distillation leveraging VQ techniques. Furthermore, ReaLM performs feature reconstruction using RVQ to learn discrete codes. Specifically, SSQR calculates the quantization loss , the structure loss , and the . Finally, these components are combined into a joint objective for optimization:
The structural loss is designed to reconstruct the KG topology (i.e., links) using the learned code representations. Meanwhile, the semantic loss distills textual information, after embedding via LLMs, into the discrete codes. In contrast, ReaLM first learns semantic KG embeddings using RotatE. The model then requires the quantized embedding, defined as , to effectively reconstruct these semantic embeddings, where represents the candidate codeword representation. Incorporating the residual quantization loss, the joint optimization objective is formulated as:
where is the residual vector.
To illustrate the integration process with LLMs, we present case studies for SSQR and ReaLM in Tables 3 and 4, respectively. These examples provide a comprehensive comparison of their distinct instruction formats and the resulting LLM outputs. As highlighted in red within the tables, there are two primary differences: SSQR prompts the model to generate quantized representations for the top-3 potential answer entities, whereas ReaLM only requires the top-1. Consequently, these differing instructions lead to distinct model outputs.
Appendix E Quantizing Graphs and Other Modalities
Current studies primarily focus on pure graph structures or text-attributed graphs Liu et al. (2025); Bo et al. (2025), which may limit their real-world applicability given the increasing importance of integrating multi-source and multi-modal information Yin et al. (2024). Following this trend, MedTok Su et al. (2025) jointly encodes textual descriptions from Electronic Health Records (EHRs) and KG-based representations of medical codes to generate richer and structured embeddings for medical foundation models. Looking forward, quantizing graphs alongside other modalities Zha et al. (2025); Ma et al. (2025), such as images or videos in general domains and electroencephalography (EEG) or electrocardiogram (ECG) in medical domains, could further enhance the cognitive capabilities of these models and extend their utility to a broader range of scenarios.
Appendix F Graph Quantization Analysis
Although several recent approaches leverage QGR to map nodes or subgraphs to discrete tokens. However, it remains an open question whether these quantized tokenizers effectively capture high-level, transferable graph patterns across diverse domains. Thus, the study Anonymous (2025) presents a comprehensive empirical study analyzing the representational consistency of quantized graph tokens across different datasets. The Graph Token Information Discrepancy (GTID) score is introduced to quantify how well structural and feature information aligns between source and target graphs for a given token. The findings reveal that current quantized tokenizers frequently assign identical tokens to structurally inconsistent patterns, leading to high GTID scores and diminished transfer performance.